Two rainstorms occurred in one week in a certain area. The first storm lasted 15 hours, and the second storm lasted 25 hours, for a total 1250mL of rain. What was the rate of rainfall for each of the two storms if the sum of the two rates was 60mL per hour?

Answer by munisa(5)

Let the rate of rainfall for the 1st rain be x ml/hr and for the 2nd rain be y ml/hr.
Volume of rain during first storm=x*15
Volume of rain during second storm=y*25
Total volume=1250ml
1250 =15x + 25y


Sum of two rates=60
x + y=60
x =60 - y
Substitute the value of x in the first equation:
1250=15(60-y) +25y
1250=900-15y+25y
350 =10y
y =35
x=60-35=25